the density of a material in C.G.S system is 8 g/cm cube. In a system of units in which unit of length is 5 cm and of mass is 20 g, the density of the material is?

Grade:12th pass

1 Answers

Agrim
6 years ago
In cgs system unit is in g/cm cubeA/QUnit of mass is gramUnit of length is cmDensity=mass/volumeV=length *breadth*height =cm*cm*cm =cm cubeD=m/v =g/cm cubeThis means that this system of units is the same as cgs systemTherefore,the density of material is same as cgs systemD= 8 g/cm cube